Acook buys two identical bags ofrice and uses some of the rice in each bag
so that one bag is half full and the other is one-third full. The cook combines them
into one bag, which then contains 33 cups ofrice. Howmuch rice was in a fullbag?

Answer by macston(5194)

.
x=amount in full bag
.
(1/3)x+(1/2)x=33 cups
(2/6)x+(3/6)x=33 cups
(5/6)x=33cups
5x=198 cups
x=39.6 cups
.
ANSWER: A full bag contains 39.6 cups.